Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name Z.E. Meziani
Log entry time 08:19:16 on November12,2008
Entry number 248374
keyword=Shift Summary
12:00 am Start of shift
Beam not in yet.
01:05 MCC called: turned the raster off and asked for the BeO target.
01:20 Beam is at BPMA (1.5,1.1) and BPMB (0.0,1.0)
01:40 Harp scans performed and logged sigmas < 200um
02:20 MCC is setting up Hall B and because Hall B slits are wide open they cannot send beam in Hall A.
02:25 MCC is ready to send beam in Hall A
asked for 6x6 for the raster. Spot++ gives 5x4uA
02:35 put the C foils in and check for the position of the hole with spot++. See log, the spot is at most 0.5mm above center rather the projected 2mm.
03:05 Move the target to reference cell with vacuum and asked for 2x2 mm raster.
run# 3265 to check spot ++
03:10 started a series of studies with raster size changes and BPMB x position changes. See elog for summary
04:35 Beam position of choice on BPMA (1.,1.3) on BPMB (-1.5,1.)
Raster size 3x3 mm and target reference cell with 100psi N2
Run # 3279, N2 at 31 psi to check beam position with spot++
Run # 3281, N2 at 100 psi 1M evts
Run # 3282, N2 at 100 psi 1M evts
6:10 Vent the N2 and fill hith H2.
Run # 3284 H2 at 100 psi dead time is high wrong prescale factor
Run # 3285 H2 at 100 psi 1M evts
Run # 3286 H2 at 100 psi 1M evts
7:40 Start beam tuning with 2x2 mm and 1 uA on pol-3He target.
BPMA (1,1.3) BPMB (-1.5,1.04)
7:50 Follow up with 3x3mm and 1uA on pol-3He good spot++
8:05 tuning with raster 4x4 mm and 1 uA good spot++
8:13 Decided to run with 3x3mm which correspond to (3x4 MCC)
for production and beam BPMA(1,1.2) and BPMB (-1.5,1)
Just start production
End of shift